a two digit number is such that the product of its digit it equal to 20.if 9 is added to the number the digits interchange their places.find the number

Answer by malaydassharma(59)

let x= once digit and y is tenth digit.
Then , number is 10y+x
So, if the digits interchange their places, then the number is 10x+y
now,
xy=20 ....(i)
10y+x+9=10x+y
or 9x-9y=9
or x-y=1 ..(ii)
From (i) and (ii) , we get
x(x-1)=20=4X5
x=4 or 5 , but as x-y=1 , then x=5 and y=4
The number is 45.